The transport of Na+ ions across a membrane, from a region where the concentration is 1 μM to a region where it is 100 μM, is coupled to the hydrolysis of ATP. Transport is opposed by a membrane potential of 70 mV. Assuming energy coupling is 100% efficient, the hydrolysis of 1 mole of ATP can drive the transport of how many moles of Na+ at 37˚C if the ΔG for ATP hydrolysis is -53.4 kJ/mol under intracellular conditions?
Answer #1

△G = RT ln([Na+]in)/([Na+]out) + ZF△⍦

R = 8.3145 J mol-1

T = 37+273 = 310K

Z = 1 for Na+

△G =(8.3145)(310 K)(ln[1/100]) + (1) (96485J*V*mol)(-0.07V)

△G = 18.62kJ/mol

1 mole ATP can give -53.4 kJ/mol

It can drive the transport of 53.4/18.62 = 2.86 approximately 3 moles of Na+